OpenStudy (anonymous):

there are 3500 cartons of apples loading in to a ship with 18 compartments. large compartment hold 250 cartons and small compartment holds 150 cartons. how many cartons will be placed in each compartment? i need help with the working out pliz?

OpenStudy (jack1):

a + b =18 250a + 150b = 3500 a = large compartment b = small compartment

OpenStudy (jack1):

do you follow where I got those equations from so far? If not I can go back and explain further?

OpenStudy (jack1):

sorry, those should be plural : a = (number of) large compartments b = (number of) small compartments

OpenStudy (jack1):

you still there champ?

OpenStudy (jack1):

assuming you're still viewing this, rearrange the first equation to be: b = 18 - a

OpenStudy (jack1):

sub that into 2nd equation: 250a + 150b = 3500 250a + 150(18 - a) = 3500 250a + 2700 - 150a = 3500 250a - 150a = 3500 - 2700 100a = 800 a = 8

OpenStudy (jack1):

therefore if: a + b =18 8 + b = 18 b = 10

OpenStudy (jack1):

so the ship has 8 large compartments and 10 small

OpenStudy (anonymous):

Good job @Jack1

OpenStudy (anonymous):

thanks very much Jack for your help cheers